# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v4

@@ Appendix C Generalized 2D bilayer mapping
Under the 2D mapping, the uniform hopping coupling in the Hamiltonian is mapped to the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling in the $x$-direction
@@
Appendix D Semiclassical approximation of the critical states
@@
Figure 7: Phase space trajectories for extended, localized and critical states. The extended state is continuous in $x$ direction while the localized state is continuous along momentum direction. The critical state is continuous in both directions. Here the extended, localized and critical states are realized in the extended AA model by setting $(t,V,\mu)=(1,2.5,0)$, $(t,V,\mu)=(1,0.5,0)$ and $(t,V,\mu)=(1,2.5,1.5)$, respectively.
@@
$E(p,x)=[2t+2\mu\cos(2\pi\alpha x)]\cos p+V\cos(2\pi\alpha x).$ (84)
Here $x=j$ labels the site coordinate and $p$ is the momentum, so that the contours $E(p,x)=E$ define the phase-space trajectories Albert and Leboeuf (2010); Malla and Raikh (2018) of the extended AAH model. As depicted in Fig. 7, the extended and localized states correspond to the trajectories continuous in $x$ direction but discontinuous in $p$ direction, whereas localized states exhibit the opposite behavior. The critical states are continuous in both directions. The dual transformation here is swapping $2\pi x$ and $p$. One can notice that $2\mu\cos(2\pi\alpha x)\cos p$ remains invariant under the dual transformation and always percolates in both directions, responsible for the critical states.
@@
$h_{0}$ $=V_{0}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{0}\cos p+2\mu_{0}\cos\bigl(2\pi\alpha x\bigr{missing})\cos p,$ (86)
$h_{x}$ $=\lambda_{x}+V_{x}\cos\bigl(2\pi\alpha x\bigr{missing})+2|t_{x}|\cos\bigl(p+\phi_{t_{x}}\bigr{missing})$
$+2|\mu_{x}|\cos\bigl(2\pi\alpha x\bigr{missing})\cos\bigl(p+\phi_{\mu_{x}}\bigr{missing}),$ (87)
$h_{y}$ $=\lambda_{y}+V_{y}\cos\bigl(2\pi\alpha x\bigr{missing})+2|t_{y}|\cos\bigl(p+\phi_{t_{y}}\bigr{missing})$
$+2|\mu_{y}|\cos\bigl(2\pi\alpha x\bigr{missing})\cos\bigl(p+\phi_{\mu_{y}}\bigr{missing}),$ (88)
$h_{z}$ $=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{z}\cos p$
$+2\mu_{z}\cos\bigl(2\pi\alpha x\bigr{missing})\cos p.$ (89)
@@
$h_{z}=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{z}\cos p,$ (90)
@@
Appendix E Exact solvability of QP systems from local constraint
Now we show the exact solvability for the spinful quasiperiodic system when the hopping coupling matrices $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate
$\det|\Pi_{j}|=0\quad{\rm or}\quad\det|\Pi_{n}|=0,$ (91)
with off-diagonal terms of $M_{j}$ being purely constant or purely quasiperiodic.
We first consider the constraint for spin-conserved process, in which case the hopping coupling matrix after the local unitary transformation is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}1&0\\ 0&0\end{array}\right).$ (92)
Without losing generality, we set $t=1$ here, then the eigenvalue equations $H\lvert\Psi\rangle=E\lvert\Psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j-1\uparrow}+u_{j+1\uparrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (93)
Notice that here $u_{j\uparrow}$ and $u_{j\downarrow}$ refer to the coefficients of the wavefunction of the dressed particle after the unitary transformation. We can then obtain the reduced eigenvalue equation
$u_{j-1\uparrow}+u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{|M_{j}^{12}|^{2}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow},$ (94)
which describe the effective motion with nearest-neighbor hopping and the effective on-site potential is energy dependently dressed by the local modulation from the accompanied internal degree of freedom.
From this expression, we can also see the reason why we require the off-diagonal terms of $M_{j}$ being purely QP or purely constant. If the off-diagonal terms of $M_{j}$ contain a constant plus a QP term, then its square will contain both a single QP frequency and double QP frequency, which generically is not analytically solvable.
We then investigate the constraint for spin-conserved process, in which case the hopping coupling matrix after the local unitary transformation is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}0&1\\ 0&0\end{array}\right),$ (95)
Again we we set $\mu=1$ here for simplicity, then the eigenvalue equations $H\lvert\psi\rangle=E\lvert\psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j-1\downarrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ u_{j+1\uparrow}+M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (96)
Then the effective eigen-equation for one for the species is
$\frac{M_{j-1}^{21}}{E-M_{j-1}^{22}}u_{j-1\uparrow}+\frac{M_{j}^{12}}{E-M_{j}^{22}}u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{1}{E-M_{j-1}^{22}}+\frac{|M_{j}^{12}|^{2}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow},$ (97)
which describes the the effective hopping and on-site potentials are energy dependently dressed by the transfer and local modulation of the accompanied internal degree of freedom.
For the combination of both are similar, where the hopping coupling matrix is given by
$\tilde{\Pi}_{j}=\left(\begin{array}[]{cc}1&1\\ 0&0\end{array}\right),$ (98)
Similarly we we set both $t=1$ and $\mu=1$, then the eigenvalue equations $H\lvert\psi\rangle=E\lvert\psi\rangle$ with generic on-site coupling matrix $M_{j}$ become
$\begin{split}u_{j+1\uparrow}+u_{j+1\downarrow}+u_{j-1\uparrow}+M_{j}^{11}u_{j\uparrow}+M_{j}^{12}u_{j\downarrow}=Eu_{j\uparrow},\\ u_{j-1\uparrow}+M_{j}^{21}u_{j\uparrow}+M_{j}^{22}u_{j\downarrow}=Eu_{j\downarrow}.\end{split}$ (99)
And the effective eigen-equation becomes
$\Big{(}1+\frac{M_{j}^{12}}{E-M_{j}^{22}}\Big{)}u_{j-1\uparrow}+\Big{(}1+\frac{M_{j+1}^{21}}{E-M_{j+1}^{22}}\Big{)}u_{j+1\uparrow}+\Big{(}M_{j}^{11}+\frac{1}{E-M_{j+1}^{22}}+\frac{M_{j}^{12}M_{j}^{21}}{E-M_{j}^{22}}\Big{)}u_{j\uparrow}=Eu_{j\uparrow}.$ (100)
Appendix F 1D QP mosaic model
In this section, we introduce the one-dimensional quasiperiodic mosaic model Wang et al. (2020a); Zhou et al. (2023) and rewrite it into our 1D spinful QP framework. The generic QP mosaic lattice refers that the modulation of on-site potential and hopping coefficients are in a mosaic manner, namely the the QP modulation or the uniform modulation only appears every $\kappa$ sites, as we will see in the following.
where the on-site potential is given by
$V_{j}=\begin{cases}V_{0}\cos(2\pi\alpha j+\theta),&j=0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\\ 0,&j\neq 0\quad{\rm mod}\leavevmode\nobreak\ \kappa.\end{cases}$ (102)
and the on-site potential given by
Extended Critical Localized
Type I $E<|\lambda^{2}/V_{0}|$ none $E>|\lambda^{2}/V_{0}|$
Type II none $E<|\lambda|$ $E>|\lambda|$
$V_{j}^{\mathrm{M,II}}$ $=$ $2tV_{j}^{\mathrm{d}}(\sigma_{0}+\sigma_{x}).$ (108)
Figure 8: The proposed laser configuration enables the realization of a model featuring all three fundamental types of mobility edges. Laser beams ${\bf E}_{1}$ and ${\bf E}_{2}$ with wavelengths $\lambda_{1}$ and $\lambda_{2}$ ,individually generate spin-independent lattices, while ${\bf E}_{3}$ and ${\bf E}_{4}$ operating at ‘tune-out’ wavelength $\lambda_{3}$ form a periodic Raman potential $(\,^{87}\text{Rb atoms})$.
Appendix G Experimental Scheme
In this section, we provide a more detailed description of the experimental scheme for realizing the Hamiltonian Eq. (57) in the main text. Without loss of generality, we consider ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$ and ||$\ket{\downarrow}\equiv\ket{1,0}$ of ${}^{87}{\rm Rb}$ atoms as an example, while all results remain applicable to other alkali atoms. After applying the unitary transformation $\sigma_{z}\rightarrow\sigma_{x}$ and $\sigma_{x}\rightarrow-\sigma_{z}$, the target Hamiltonian can be rewritten as
$H=\left[\frac{p_{z}^{2}}{2m}+{\cal V}_{s}(z)\right]\otimes\sigma_{0}+{\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2-M_{0}\sigma_{z},$ (109)
@@ Appendix G Experimental Scheme
${\cal V}_{1}(z)=\frac{V_{1}}{2}\cos(2k_{1}z-\phi_{1}^{\prime}),$ (111)
where
$V_{1}=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,10}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (112)
In the above, the transition matrix elements are $\Omega_{F^{\prime},m_{F}^{\prime};\sigma,nq}^{(J)}=-\langle F^{\prime},m_{F}^{\prime}|{\bf d}|\sigma\rangle E_{n}/\hbar$ and $\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}$ are detunings of laser field from different levels. Here, $J=1/2$, $3/2$ correspond to the $D_{1}$ and $D_{2}$ lines respectively, and $\sigma=\uparrow,\downarrow$ labels the spin states. The index $n$ denotes the laser beam, while $q=0,\pm$ represents the polarization of laser field. For a laser wavelength of $\lambda_{1}\approx 1064\,{\rm nm}$, the overall optical shifts induced by the linearly polarized ($q=0$) field from ${\bf E}_{1}$ are nearly identical for the two spin states. Comparing this with Eq. (57), we find
$k_{s}=2k_{1},\ V_{s}=V_{1}/2,\ \phi_{s}=-\phi_{1}^{\prime}.$ (113)
On the other hand, the primary lattice ${\cal V}_{p}(z)$ is introduced through three laser fields, given by
@@
${\cal V}_{p}(z)=\frac{V_{2}}{2}\cos(2k_{2}z-\phi_{2}^{\prime}),$ (115)
where
$V_{2}=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,20}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (116)

### v5

@@ Appendix D Generalized 2D bilayer mapping
Under the 2D mapping, the uniform hopping in Hamiltonian is mapped to the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling along $x$-direction
@@
Appendix E Exact solvability of QP systems from local constraint
In this section, we provide the details of the exact solvability of the spinful QP lattice model under the local constraint $\det|\Pi_{j}|=0$ or $\det|\Pi_{n}|=0$.
We begin by briefly reviewing Avila’s global theory for one-frequency QP Schrödinger operators Avila (2015). Let $T$ be an analytic function mapping the circle $S^{1}$ into the group $SL(2,\mathbb{C})$. An analytic cocycle $(\alpha,T)$ is defined as a linear skew product
$(\alpha,T):S^{1}\times\mathbb{R}^{2}\rightarrow S^{1}\times\mathbb{R}^{2},\qquad(\theta,\psi)\mapsto(\theta+\alpha,T(\theta)\,\psi).$
[…]
$\gamma_{\epsilon}(T)=\lim_{n\to\infty}\frac{1}{n}\int\ln\|T_{n}(\theta+i\epsilon)\|\,d\theta,$ (88)
where $T_{n}$ denotes the transfer matrix at site $n$.
Imposing the exact-solvability condition, the original spinful QP lattice model can be rigorously reduced to an effective 1D tight-binding problem with NN hopping $t^{\rm eff}_{j}$ and onsite potential $V^{\rm eff}_{j}$. The resulting transfer matrix inherits a constrained analytic structure, allowing the LE to be evaluated explicitly via Avila’s global theory.
E.1 New exactly solvable mosaic models
We present the analytic derivation for the LE of the QP spin-flipped (QPSF) model in Eq. (21). The effective hopping and onsite potential are given by
$t_{j}^{\rm eff}=2\lambda t\cos(2\pi\alpha j),\quad V_{j}^{\rm eff}=\lambda^{2}+4t^{2}\cos^{2}(2\pi\alpha j).$
The corresponding complexified transfer matrix can be written in the factorized form $T_{j}(\theta+i\epsilon)=\tilde{T}_{j}(\theta+i\epsilon)/(2\cos(\theta+\alpha/2+i\epsilon))$, where the numerator matrix $\tilde{T}_{j}$ has elements $\tilde{T}_{j}^{11}=[E^{2}-\lambda^{2}-4t^{2}\cos^{2}(\theta+i\epsilon)]/\lambda t$, $\tilde{T}_{j}^{12}=-2\cos(\theta-\alpha+i\epsilon)$, $\tilde{T}_{j}^{21}=2\cos(\theta+i\epsilon)$, and $\tilde{T}_{j}^{22}=0$. The explicit factor $2\cos(\theta)$, which corresponds to the IDZs discussed in the main text, immediately excludes absolutely continuous spectral components and therefore forbids extended states. Taking the asymptotic limit $\epsilon\to\infty$, one finds $\tilde{T}_{j}=e^{-i\theta+2\epsilon}[T^{\prime}+\mathcal{O}(e^{-\epsilon})]$, where the only nonvanishing element of $T^{\prime}$ is $T^{\prime}_{11}=-t/\lambda$.
Now we calculate the LE. Applying Jensen’s formula, we obtain
$\int_{0}^{2\pi}\ln|2\cos(\theta+i\epsilon)|{\mathrm{d}\theta}=2\pi|\epsilon|.$ (89)
The total complexified LE is therefore
$\begin{split}2\gamma_{\epsilon}=&\lim_{m\to\infty}\frac{1}{2\pi m}\int_{0}^{2\pi}\ln\left\|\tilde{T}_{j+m-1}\cdots\tilde{T}_{j}\right\|\mathrm{d}\theta\\ &-\int_{0}^{2\pi}\ln|2\cos(\theta+\alpha/2+i\epsilon)|{\mathrm{d}\theta}\\ =&\ln\|T^{\prime}\|+2|\epsilon|-|\epsilon|+\mathcal{O}\left(\epsilon^{-1}\right)\\ =&\ln|t/\lambda|+|\epsilon|+\mathcal{O}\left(\epsilon^{-1}\right)\\ \end{split}.$
Since $2\gamma_{\epsilon}(E)$ is a convex, piecewise-linear function of $\epsilon$ with quantized slopes, this asymptotic behavior uniquely fixes the Lyapunov exponent for all $\epsilon\in\mathbb{R}$ as $2\gamma_{\epsilon}=\max\left\{0,\ln|t/\lambda|+|\epsilon|\right\}$. Setting $\epsilon=0$, we obtain
$\gamma=\max\left\{0,\frac{1}{2}\ln|t/\lambda|\right\},$ (90)
Which is the LE in the main text. A positive LE corresponds to the localized phase, and for the zero LE, it corresponds to the critical states. Finally, applying the duality transformation yields the phase diagram of the corresponding dual mosaic model in the main text.
E.2 The spin-selective QP lattice model
We procede to give the details of the SSQP model. As discussed in the main text, the SSQP model and its dualcounterpart are exactly solvable at the condition $t=V_{0}=V_{B}=0$ and $\lambda_{0}=\lambda_{1}=\lambda$, for which the effective hopping and onsite potential reduce to the forms given in Eq. (29). Following the same strategy as above, we decompose the complexified transfer matrix as $T_{j}(\theta+i\epsilon)=\tilde{T}_{j}(\theta+i\epsilon)/(2\cos(\theta+\alpha/2+i\epsilon))$, with
$\tilde{T}_{j}=\begin{pmatrix}\frac{E}{\mu}-\frac{\lambda^{2}(4+4\cos(\theta+i\epsilon))}{\mu E}&-2\cos(\theta-\alpha/2+i\epsilon)\\ 2\cos(\theta+\alpha/2+i\epsilon)&0\end{pmatrix}$,
Taking the limit $\epsilon\to\infty$, we obtain $\tilde{T}_{j}=e^{-i\theta+\epsilon}[T^{\prime}+\mathcal{O}(e^{-\epsilon})]$, with the $T^{\prime}$ given by
$T^{\prime}=\begin{pmatrix}{2\lambda^{2}}/{\mu E}&-e^{i\alpha/2}\\ e^{-i\alpha/2}&0\end{pmatrix}$ (91)
Combining with Jenson’s formula, the LE takes the form
[…]
Since $2\gamma\epsilon(E)$ is a convex, piecewise-linear function of $\epsilon$ with quantized slopes, this asymptotic behavior uniquely determines
$\gamma_{\epsilon}(E)=\frac{1}{2}\max\left\{0,\ln|\big|\lambda^{2}/\mu E\big|+\sqrt{(\lambda^{2}/\mu E)^{2}-1}\Big|\right\}.$ (92)
Setting $\epsilon=0$ recovers the Lyapunov exponent quoted in Eq. (30) of the main text.
We next provide analytic details for the SSQP model in the case $\mu=\lambda_{1}=0$, where the effective hopping and onsite potential are given in Eq. (32). The transfer matrix can be factorized as $T_{j}=T_{j}^{A}T_{j}^{B}$, with $T_{j}^{A}=t/(E-V_{0}-2V_{B}V_{j}^{d})$ and
$T_{j}^{B}=\begin{pmatrix}\frac{E(E-V_{0}-2V_{B}V_{j}^{d})}{t^{2}}-\lambda_{0}^{2}&-\frac{E-V_{0}-2V_{B}V_{j}^{d}}{t}\\ \frac{E-V_{0}-2V_{B}V_{j}^{d}}{t}&0\end{pmatrix}$,
Accordingly, the LE decomposes as $2\gamma=\gamma_{A}+\gamma_{B}$.To evaluate $\gamma_{B}$, we complexify $\theta\to\theta+i\epsilon$ and take $\epsilon\to\infty$, yielding $T_{j}^{B}=e^{-i\theta+\epsilon}[\mathcal{T}^{B}+\mathcal{O}(e^{-\epsilon})]$, with $\mathcal{T}^{B}_{11}=-{EV_{B}}/{t^{2}}$, $\mathcal{T}^{B}_{12}={V_{B}}/{t}$, $\mathcal{T}^{B}_{21}=-{V_{B}}/{t}$, and $\mathcal{T}^{B}_{22}=0$. And the complexified $\gamma_{B}$ up to order $\mathcal{O}(\epsilon^{-1})$ is given by
$\gamma_{B}=\ln|\frac{|EV_{B}/t^{2}|+\sqrt{(EV_{B}/t^{2})^{2}-(2V_{B}/t)^{2}}}{2}\Big|+2\pi|\epsilon|.$
Now we calculate $\gamma_{A}$ under different parameter regimes. When $|E-V_{0}|>2V_{B}$, one finds
$\begin{split}\gamma_{A}=&\int_{0}^{2\pi}\ln|\frac{t}{E-V_{0}-2V_{B}\cos(\theta+i\epsilon)}\Big|d\theta\\ =&\ln|\frac{t}{V_{B}}\Big|-2\pi|\epsilon|,\end{split}$
Therefore we have the total Lyapunov exponent given by
$\gamma(E)=\ln|\frac{|E/t|+\sqrt{(E/t)^{2}-4}}{2}\Big|,$ (93)
which gives the LE in the main text Eq. (33). When $|E-V_{0}|<2V_{B}$, the $\gamma_{A}$ is given by
$\begin{split}\gamma_{A}=&\int_{0}^{2\pi}\ln|\frac{t}{E-V_{0}-2V_{B}\cos(\theta+i\epsilon)}\Big|d\theta\\ =&\ln\left|\frac{2}{|\chi_{0}|+\sqrt{\chi_{0}^{2}-\chi_{B}^{2}}}\right|,\end{split}$
and the total Lyapunov exponent becomes
$\gamma(E)=\ln|\frac{|\chi_{E}|+\sqrt{\chi_{E}^{2}-\chi_{B}^{2}}}{|\chi_{0}|+\sqrt{\chi_{0}^{2}-\chi_{B}^{2}}}\Bigg|+2\pi|\epsilon|,$ (94)
where $\chi_{B}=2V_{B}/t$, $\chi_{E}=EV_{B}/t^{2}$ and $\chi_{0}=(E-V_{0})/t$. Setting $\epsilon=0$ reproduces Eq. (34) of the main text.
Supplementary Material
Appendix S-1 Semiclassical approximation of the critical states
@@ E.2 The spin-selective QP lattice model
Figure S1: Phase space trajectories for extended, localized and critical states. The extended state is continuous in $x$ direction while the localized state is continuous along momentum direction. The critical state is continuous in both directions. Here the extended, localized and critical states are realized in the extended AA model by setting $(t,V,\mu)=(1,2.5,0)$, $(t,V,\mu)=(1,0.5,0)$ and $(t,V,\mu)=(1,2.5,1.5)$, respectively.
@@
$E(p,x)=[2t+2\mu\cos(2\pi\alpha x)]\cos p+V\cos(2\pi\alpha x).$ (S1)
Here $x=j$ labels the site coordinate and $p$ is the momentum, so that the contours $E(p,x)=E$ define the phase-space trajectories Albert and Leboeuf (2010); Malla and Raikh (2018) of the extended AAH model. As depicted in Fig. S1, the extended and localized states correspond to the trajectories continuous in $x$ direction but discontinuous in $p$ direction, whereas localized states exhibit the opposite behavior. The critical states are continuous in both directions. The dual transformation here is swapping $2\pi x$ and $p$. One can notice that $2\mu\cos(2\pi\alpha x)\cos p$ remains invariant under the dual transformation and always percolates in both directions, responsible for the critical states.
@@
$h_{0}$ $=V_{0}\cos\bigl(2\pi\alpha x\bigr.)+2t_{0}\cos p+2\mu_{0}\cos\bigl(2\pi\alpha x\bigr.)\cos p,$ (S3)
[…]
$h_{z}$ $=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr.)+2t_{z}\cos p$
$+2\mu_{z}\cos\bigl(2\pi\alpha x\bigr.)\cos p.$ (S6)
@@
$h_{z}=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr.)+2t_{z}\cos p,$ (S7)
where $t_{z}$ is the uniform hopping coefficient, and $V_{z}$ is the QP on-site coupling, contributing IDZs in $\sigma_{z}$ component. The $h_{z}^{2}$ in Eq. (S2) produces the cross term $V_{z}\,2t_{z}\cos\bigl(2\pi\alpha x\bigr.)\cos p$, which gives rise to the critical orbitals in the system. One can see that the IDZs in the matrix component of $M_{j}$ shared in $\Pi_{j}$ leads to the percolated phase-space trajectory, which is a mechanism unique to spinful QP systems.
Appendix S-2 Evolution of phase boundaries away from exactly solvable limits
Upon departing from the exactly solvable limits, the phase boundaries generically evolve smoothly in parameter space; however, whether the spectrum develops a complicated structure of mobility edges (MEs) depends crucially on the symmetries preserved in the corresponding parameter regime. In particular, chiral symmetry can forbid MEs altogether, leading to smooth deformations of phase boundaries, whereas broken chiral symmetry allows energy-dependent restructuring of the spectrum and the emergence of intricate ME structures.
In this section, we show that upon departing from the exactly solvable limits, the phase boundaries generically evolve smoothly in parameter space; however, whether the spectrum develops a complicated structure of mobility edges (MEs) depends crucially on the symmetries preserved in the corresponding parameter regime. In particular, chiral symmetry can forbid MEs altogether, leading to smooth deformations of phase boundaries, whereas broken chiral symmetry allows energy-dependent restructuring of the spectrum and the emergence of intricate ME structures.
When the chiral symmetry is preserved, MEs are absent and the phase boundaries deform smoothly upon departing from the exactly solvable point. This behavior is illustrated in Fig. S2, where we introduce an imbalance between $t_{0}$ and $t_{\rm so}$ while keeping $\eta=1$. In this case, the phase boundaries between critical-localized and critical-extended phases shift continuously according to $M_{z}=2|t_{0}\pm t_{\rm so}|$.
Figure S2: Departing from exactly solvable limits. Phase diagrams at fixed $\eta=1$ illustrating smooth deformations away from the exactly solvable point. The fractal dimension (FD) as functions of $M_{z}/t_{0}$ and energy $E/t_{0}$ for $t_{\rm so}/t_{0}=1.0$ (a), $0.9$ (b), and $0.8$ (c), respectively. The dashed vertical lines mark the analytically predicted phase boundaries $M_{z}=2|t_{0}\pm t_{\rm so}|$. Despite the imbalance between $t_{0}$ and $t_{\rm so}$, the absence of mobility edges persists due to the chiral(-like) symmetry, and the phase boundaries shift continuously.
Figure S3: Departing from exactly solvable limits. Phase diagrams at fixed $\eta=0$ illustrating the emergence of mobility edges away from the exactly solvable point. The fractal dimension (FD) as functions of $M_{z}/t_{0}$ and energy $E/t_{0}$ for $t_{\rm so}/t_{0}=1.0$ (a), $0.9$ (b), and $0.8$ (c), respectively. Upon introducing an imbalance between $t_{0}$ and $t_{\rm so}$, extended states appear at small $M_{z}$ and energy-dependent mobility edges develop.
Figure S4: Departing from exactly solvable limits. Phase diagrams at fixed $\eta=0.5$ showing the development of mobility edges. The fractal dimension (FD) as functions of $M_{z}/t_{0}$ and energy $E/t_{0}$ for $t_{\rm so}/t_{0}=1.0$ (a), $0.9$ (b), and $0.8$ (c), respectively. In contrast to the chiral-symmetric case, the imbalance between $t_{0}$ and $t_{\rm so}$ induces energy-dependent effective hopping processes (see main text), leading to extended states and nontrivial mobility edges.
Figure S5: Departing from exactly solvable limits. Phase diagrams illustrating the evolution of mobility edge upon varying $\eta$ at fixed $t_{\rm so}=t_{0}$. The fractal dimension (FD) as functions of $M_{z}/t_{0}$ and energy $E/t_{0}$ for $\eta=0.2$ (a), $0.4$ (b), $0.6$ (c), and $0.8$ (d), respectively. Away from the exactly solvable points $\eta=0,\,0.5,\,1$, $\eta\neq 1$ breaks chiral symmetry and introduces energy dependent quasiperiodic hopping processes, leading to critical states and associated mobility edges.
In contrast, when the chiral symmetry is broken, the phase diagram can develop a more intricate structure of MEs. Specifically, for $\eta=0$ and $\eta=0.5$, as shown in Figs. S3 and S4, relaxing the condition $t_{0}=t_{\rm so}$ leads to the MEs associated with extended states. This behavior can be attributed to the effective next-nearest-neighbor hopping induced by the imbalance between $t_{0}$ and $t_{\rm so}$, which has an energy-dependent form
$t_{j}^{\rm NN}=\frac{t_{+}t_{-}}{E-\Delta_{j+1}},$ (S8)
where $t_{\pm}=t_{0}\pm t_{\rm so}$. This process favors extended states at small $M_{z}$ but depends nontrivially on energy, thereby producing complex ME structures. At sufficiently large $M_{z}$, these additional processes become subdominant and the overall phase diagram remains qualitatively unchanged.
A similarly rich ME structure also arises when departing from exact solvability by varying $\eta$ while keeping $t_{\rm so}=t_{0}$. In this case, the model lacks chiral symmetry, and a finite $\eta$ simultaneously introduces energy dependent IDZs and higher-order (quartic) contributions in $M_{z}$, leading to energy-dependent localization transitions. As shown in Fig. S5, varying $\eta$ from 0 to 1 interpolates between the three exactly solvable points $\eta=0,0.5,1$, during which complicated MEs generically emerge. This behavior can be understood from the effective quasiperiodic hopping
$t_{j}^{\rm eff}=\frac{-2t_{0}\eta\Delta_{j}}{E-(1-\eta\Delta_{j})},$ (S9)
which induces critical states, together with the effective onsite potential
$V_{j}^{\mathrm{eff}}=\frac{4t_{0}^{2}}{E-(1-\eta)\Delta_{j-1}}+\frac{E(1-\eta)\Delta_{j}-(1-2\eta)\Delta_{j}^{2}}{E-(1-\eta)\Delta_{j}}.$ (S10)
For $0<\eta<0.5$ and $0.5<\eta<1$, the coexistence of single- and double-frequency quasiperiodic modulations renders the locations of MEs highly nontrivial.
Appendix S-3 Experimental Scheme
@@
$H=\left[\frac{p_{z}^{2}}{2m}+{\cal V}_{s}(z)\right]\otimes\sigma_{0}+{\cal V}_{p}(z)(\sigma_{0}-\sigma_{x})/2-M_{0}\sigma_{z},$ (S11)
@@
[…]
Figure S6: The proposed laser configuration enables the realization of a model featuring all three fundamental types of mobility edges. Laser beams ${\bf E}_{1}$ and ${\bf E}_{2}$ with wavelengths $\lambda_{1}$ and $\lambda_{2}$ ,individually generate spin-independent lattices, while ${\bf E}_{3}$ and ${\bf E}_{4}$ operating at ‘tune-out’ wavelength $\lambda_{3}$ form a periodic Raman potential $(\,^{87}\text{Rb atoms})$.
In the above, the transition matrix elements are $\Omega_{F^{\prime},m_{F}^{\prime};\sigma,nq}^{(J)}=-\langle F^{\prime},m_{F}^{\prime}|{\bf d}|\sigma\rangle E_{n}/\hbar$ and $\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}$ are detunings of laser field from different levels. Here, $J=1/2$, $3/2$ correspond to the $D_{1}$ and $D_{2}$ lines respectively, and $\sigma=\uparrow,\downarrow$ labels the spin states. The index $n$ denotes the laser beam, while $q=0,\pm$ represents the polarization of laser field. For a laser wavelength of $\lambda_{1}\approx 1064\,{\rm nm}$, the overall optical shifts induced by the linearly polarized ($q=0$) field from ${\bf E}_{1}$ are nearly identical for the two spin states. Comparing this with Eq. (52), we find $k_{s}=2k_{1},\ V_{s}=V_{1}/2$ and $\phi_{s}=-\phi_{1}^{\prime}$. On the other hand, the primary lattice ${\cal V}_{p}(z)$ is introduced through three laser fields, given by
@@
[…]